The SP-6 Index (a fictitious index) is used by many investors to monitor the general behavior of the stock market. It has a base value set equal to 100 at January 1, 1978 In the table shown here.囲, the closing market values for each of the 6 stocks included in the index are given for 3 dates a. Calculate the value of the SP-6 Index on both January 1, 2016, and June 30, 2016, using the data presented here b. Compare the values of the SP-6 Index calculated in part a and relate them to the base index value. Would you describe the general market condition during the 6-month period January 1 to June 30, 2016, as a bull or a bear market? a. The value of the SP-6 Index on January 1, 2016 is(Round to two decimal places.) Data Table Closing Market Value of Stock (SThousands) January 1, 2016 Stock June 30, 2016 $431 $1,156 $976 $364 $642 S287 $469 $1,123 $984 $415 $705 S329 January 1, 1978 $230 $622 $458 $141 S313 $74 6 (Click on the icon located on the top-right corner of the data table below in order to copy its contents into a spreadsheet.) Enter vour answe
Answer #1

Answer:

A)

Closing Market Valuation
Stock June 30,2016 January 1,2016 January 1,1978
1 431 469 230
2 1156 1123 622
3 976 984 458
4 364 415 141
5 642 705 313
6 287 329 74
Total 3856 4025 1838

So value of SP-6 index on January 1,2016=$4025

So value of SP-6 index on June 30,2016=$3856

Since on January 1,1978 we have set base value=100

Value of SP-6 index on January 1 , 2016 = (4025/1838)*100=218.98

Since on January 1,1978 we have set base value=100

Value of SP-6 index on June 30 , 2016 = (3856/1838)*100=209.79

B)

Since value of SP-6 index from January 1,2016 to June 30, 2016 has been decreased from $4025 to $3856 Hence it is bear market.